[figure: 9.7 cm]

MAGNETIC PROPERTIES OF GdMnAl COMPOUND

Jerzy Jarosz^{∗}, Tadeusz Mydlarz^{+}

∗Institute of Physics, University of Silesia, Uniwersytecka 4, 40-007 Katowice, Poland

+International Laboratory of High Fields and Low Temperatures, Gajowicka 95, 53-529 Wrocł aw, Poland

Single phase, policrystalline samples of GdMnAl compound were obtained in the arc furnace. The samples were identified by X-ray powder diffraction. GdMnAl crystallize in the cubic MgCu_{2} - type structure with a=7.887 Å. The XPS valence band spectrum showed that Mn 3d states are localized near the Fermi level. Measurements of magnetization versus magnetic field up to 150 kOe revealed paramagnetic character of GdMnAl and the electron paramagnetic resonance investigation showed strong bottleneck effect existing in this compound.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : J. Jarosz

Address for correspondence : Institute of Physics, University of Silesia, Uniwersytecka 4, 40-007 Katowice, Poland

Email address : jjarosz@us.edu.pl